500+22x=360+28x
Please help me solve this as simply as possible, step by step?

Answer by ilana(307)

If you do the same thing to both sides of an equation, you get another equation with the same solution.
First subtract 22x from both sides.
500 + 22x - 22x = 360 + 28x - 22x
500 = 360 + 6x
Now subtract 360 from both sides.
500 - 360 = 360 + 6x - 360
140 = 6x
Now divide both sides by 6.
140/6 = 6x/6
70/3 = x
So the solution is x = 70/3, or approximately 23.3